OpenStudy (anonymous):

Find the surface area of a pyramid with a square base with dimensions 100 mm by 100 mm and height of 75 mm.

pooja195 (pooja195):

We need the pythagorean theorem to calculate s: s = sqrt[(0.5*b)^2 + h^2] = sqrt[(0.5*100)^2 + 75^2] = 90.139 mm Therefore, A = 2(100)(90.139) + (100)^2 A = 28,027.8 mm^2

OpenStudy (anonymous):

Thank you (:

pooja195 (pooja195):

np!